Grade 12th passAlgebra

Find the ratio of
15 kg to 210 g
Find the ratio of
15 kg to 210 g

Harsh Singh Chauhan
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We know that,
1 kg = 1000 g
Then,
15 kg = 15 × 1000 = 15000 g
Now we have to find the ratio,
= 15000/210
= 1500/21
= 500/7 ... [∵divide both by 3]
So, the required ratio is 500: 7.

Thanks